OpenStudy (anonymous):

A 6-sided number cube numbered 1-6 is tossed 120 times. About how many times would you expect the cube to show an even number?

OpenStudy (anonymous):

120 x 3 because there is only 3 even numbers on the cube 2,4 and 6 , so multiply 120 by 3

OpenStudy (anonymous):

so the answer is 60

OpenStudy (anonymous):

rit?????

OpenStudy (anonymous):

umi said multiply 120 by 3 so.... nope

OpenStudy (anonymous):

. . .. 360

OpenStudy (anonymous):

p(even number)=3/6=1/2 therefore, E(even number)= 1/2*120=60 u got it right merrylin

OpenStudy (anonymous):

or no divide sorry

OpenStudy (anonymous):

so yes ur correct

OpenStudy (anonymous):

ya thths wat i thougth it was lol but thx anywayz thxxx

OpenStudy (anonymous):

ok welcum hun , srry almost gave u the wrong stradegy